OpenStudy (carlif01):

you and your friend are splitting a restaurant bill it if you pay half of the bill plus $5 your friend pays half of the bill plus 25% of half the bill you and your friend pay the same total amount.How Much is half the bill?

OpenStudy (3mar):

Any ideas?

OpenStudy (carlif01):

what do you mean?

OpenStudy (carlif01):

please help i need to have this done before 4:00p.m

OpenStudy (3mar):

Can you form an expression of what they paid?

OpenStudy (carlif01):

no that is what i need help with

OpenStudy (3mar):

Ok Step by step.

OpenStudy (carlif01):

i need to make a equation and then solve it and check it all by showing my work

OpenStudy (carlif01):

please hurry i really need to have this done before 4:00

OpenStudy (3mar):

"you and your friend are splitting a restaurant bill" means that you pay as your friend pay right?

OpenStudy (carlif01):

i guess

OpenStudy (3mar):

Great!

OpenStudy (3mar):

Let' consider that the pill has a mount of x $ ok?

OpenStudy (3mar):

"you pay half of the bill plus $5" means??

OpenStudy (carlif01):

means what?

OpenStudy (3mar):

@CarliF01

OpenStudy (carlif01):

yes

OpenStudy (3mar):

"you pay half of the bill plus $5"

OpenStudy (carlif01):

did you figure out a equation?

OpenStudy (carlif01):

what? i don't understand

OpenStudy (3mar):

"you pay half of the bill plus $5" means that you have paid (x+5)$ right?

OpenStudy (carlif01):

yes

OpenStudy (carlif01):

please hurry

OpenStudy (3mar):

"your friend pays half of the bill plus 25% of half the bill" means that your friend has paid \((\frac{x}{2}+[0.25*\frac{x}{2}])\) right?

OpenStudy (carlif01):

i figured it out its x=40 then half of that thanks

OpenStudy (3mar):

By the way, I am sorry about that mistake.. you pay half of the bill plus $5": means that you have paid \((\frac{x}{2}+5)$\)

OpenStudy (3mar):

Did you get it?

OpenStudy (3mar):

@CarliF01 please hurry